OpenStudy (anonymous):

I need to write an equation for the line containing the given point and parallel to the given line then put it in y=mx+b form. (3,4); x+2y=7

myininaya (myininaya):

ok so we have x+2y=7 write this is y=mx+b form

myininaya (myininaya):

subtract x on both sides: x-x+2y=-x+7 2y=-x+7 divide both sides by 2: y=-x/2+7/2 what is the slope here?

OpenStudy (anonymous):

1/2

myininaya (myininaya):

-1/2

myininaya (myininaya):

ok all parallel lines have the same slope

myininaya (myininaya):

so our equation parallel to the line x+2y=7 (or y=-x/2+7/2) is y=-x/2+b so we need to find the y-intercept (aka b)

myininaya (myininaya):

we know a point (3,4) on this line

myininaya (myininaya):

x is 3 y is 4 so we have 4=-3/2+b (now you solve for b)

OpenStudy (anonymous):

Now I'm lost

OpenStudy (anonymous):

I'm not sure of the math.. do I divide -3/2 by 4

myininaya (myininaya):

to get b by itself what do you add to both sides

OpenStudy (anonymous):

4

myininaya (myininaya):

no! we have 4=-3/2+b we want b by itself so we need to add 3/2 on both sides

myininaya (myininaya):

4+3/2=-3/2+3/2+b 4+3/2=0+b 4+3/2=b b=4+3/2 b=4*2/2+3/2 (common denominator) b=8/2+3/2 b=11/2 so y=-x/2+11/2

OpenStudy (anonymous):

Thanks